OpenStudy (anonymous):

Simplify 15x+15 over x^2-1 A.-15 B.15/x-1 C.30/x-1 D.15

OpenStudy (mertsj):

Factor 15 out of the numerator first

OpenStudy (mertsj):

the denominator is the difference of two squares and it factors are (x-1)(x+1)

OpenStudy (mertsj):

Then reduce the fraction by cancelling out the common factor x+1

OpenStudy (anonymous):

(15x + 15)/(x^2 - 1) = 15(x + 1)/((x - 1)(x + 1)) = 15/(x-1) B